# Elevating Visual Perception in Multimodal LLMs with Visual Embedding Distillation

### v2

@@ Appendix B Enhanced Projector post Pre-Training
We compare the probing scores for the feature outputs from the projector module of the LLaVA-1.5 and VisPer-LM models between checkpoints obtained after the PT stage (proj_PT) of training. As shown in Tab. XI, the representations in models under proj_PT are better for VisPer-LM than LLaVA-1.5, indicating the predictive embedding optimization assisting in learning an improved projector before the IFT stage.
@@ Appendix C Probing Encoder-Free MLLMs
We notice that the representation quality for features from EVE-7B is relatively worse in the initial layers, owing to the patch embedding operation that outputs coarse features. However, the probing performance improves in the deeper layers, which could be attributed to the patch-alignment operation diao2024EVE . Notably, EVE-7B uses about 35M diao2024EVE training samples but still lags slightly behind LLaVA-1.5 in performance (Tab. XII), with our VisPer-LM showing better representation quality and performance, underscoring the effectiveness of our approach.
Unlike EVE-7B, features from Chameleon-7B chameleon have surprisingly low depth and seg representation quality in the deeper layers. Unlike other models, Chameleon-7B follows a steep downward trend in the deeper layers, which we attribute to its relatively low training on VQA data with a heavy focus on text-to-text and text-to-image tasks in the SFT data mixture, which is also evident from its low performance on CV-Bench. Note that during evaluation, we noticed that Chameleon cannot output option letters required to evaluate on MCQ-type CV-Bench. Therefore, we leverage Qwen-2.5-72B-Instruct qwen2 to extract the option letters corresponding to the sentence output from Chameleon to obtain the numbers for Tab. XII.
@@ Appendix D Correlation between VQA Performance and Probing Performance
Takeaways 1. As shown in Tab. XIII, we observe the highest positive correlation between CV-Bench accuracy and seg/depth probing performance. 2. Moreover, we find a high correlation between the performance on the Depth/Relation task and Count tasks with the depth and seg probing performance, respectively. 3. As expected, we find low performance correlation with gen probe scores, indicating that gen features act as a good proxy for general visual representation targets. 4. Surprisingly, we observe a negative correlation between the Distance task and all probe scores, which could be attributed to the lack of “real-world distance” estimation queries in the training dataset.
Table XIII: Correlation of CV-Bench performance to probing performance. We notice the high positive correlation of 0.98 between depth probing performance and CV-Bench accuracy. Moreover, we find that performance on the Count and Depth subtasks also highly correlates with the seg and depth probe cosine-similarity scores, respectively. Additionally, we find a low correlation between the gen probing performance and CV-Bench performance.
@@ Appendix E Evaluating Probes on Downstream Tasks
We evaluate the probes trained against the target features on the corresponding downstream target tasks, i.e., image generation, depth estimation and image segmentation. To obtain the predictions, we feed the outputs from the probes into the decoder from the corresponding target model. Specifically, we report the FID Seitzer2020FID scores on 5k images from COCO-val2017 coco for gen probes, accuracy on the DA-2K depth_anything_v2 benchmark for depth probes, and mIoU jain2023oneformer on the COCO-val2017 coco set for seg probes. We average the scores over all layers for easier comparison. As shown in Tab. XIV, probes trained for our VisPer-LM outperform those for the baseline LLaVA-1.5 liu2023improvedllava model across all probing tasks, proving the improved visual representation quality owing to our embedding optimization approach.
@@
Table XIV: Quantitative Evaluation on target probing task. Probes trained for our VisPer-LM perform significantly better as compared to the probes trained on baseline LLaVA-1.5 liu2023improvedllava .

### v3

@@ Appendix B Enhanced Projector post Pre-Training
We compare the probing scores for the feature outputs from the projector module of the LLaVA-1.5 and VisPer-LM models between checkpoints obtained after the PT stage (proj_PT) of training. As shown in Tab. VIII, the representations in models under proj_PT are better for VisPer-LM than LLaVA-1.5, indicating the predictive embedding optimization assisting in learning an improved projector before the IFT stage.
@@ Appendix C Probing Encoder-Free MLLMs
We notice that the representation quality for features from EVE-7B is relatively worse in the initial layers, owing to the patch embedding operation that outputs coarse features. However, the probing performance improves in the deeper layers, which could be attributed to the patch-alignment operation diao2024EVE . Notably, EVE-7B uses about 35M diao2024EVE training samples but still lags slightly behind LLaVA-1.5 in performance (Tab. IX), with our VisPer-LM showing better representation quality and performance, underscoring the effectiveness of our approach.
Unlike EVE-7B, features from Chameleon-7B chameleon have surprisingly low depth and seg representation quality in the deeper layers. Unlike other models, Chameleon-7B follows a steep downward trend in the deeper layers, which we attribute to its relatively low training on VQA data with a heavy focus on text-to-text and text-to-image tasks in the SFT data mixture, which is also evident from its low performance on CV-Bench. Note that during evaluation, we noticed that Chameleon cannot output option letters required to evaluate on MCQ-type CV-Bench. Therefore, we leverage Qwen-2.5-72B-Instruct qwen2 to extract the option letters corresponding to the sentence output from Chameleon to obtain the numbers for Tab. IX.
@@ Appendix D Correlation between VQA Performance and Probing Performance
Takeaways 1. As shown in Tab. X, we observe the highest positive correlation between CV-Bench accuracy and seg/depth probing performance. 2. Moreover, we find a high correlation between the performance on the Depth/Relation task and Count tasks with the depth and seg probing performance, respectively. 3. As expected, we find low performance correlation with gen probe scores, indicating that gen features act as a good proxy for general visual representation targets. 4. Surprisingly, we observe a negative correlation between the Distance task and all probe scores, which could be attributed to the lack of “real-world distance” estimation queries in the training dataset.
Table X: Correlation of CV-Bench performance to probing performance. We notice the high positive correlation of 0.98 between depth probing performance and CV-Bench accuracy. Moreover, we find that performance on the Count and Depth subtasks also highly correlates with the seg and depth probe cosine-similarity scores, respectively. Additionally, we find a low correlation between the gen probing performance and CV-Bench performance.
@@ Appendix E Evaluating Probes on Downstream Tasks
We evaluate the probes trained against the target features on the corresponding downstream target tasks, i.e., image generation, depth estimation and image segmentation. To obtain the predictions, we feed the outputs from the probes into the decoder from the corresponding target model. Specifically, we report the FID Seitzer2020FID scores on 5k images from COCO-val2017 coco for gen probes, accuracy on the DA-2K depth_anything_v2 benchmark for depth probes, and mIoU jain2023oneformer on the COCO-val2017 coco set for seg probes. We average the scores over all layers for easier comparison. As shown in Tab. XI, probes trained for our VisPer-LM outperform those for the baseline LLaVA-1.5 liu2023improvedllava model across all probing tasks, proving the improved visual representation quality owing to our embedding optimization approach.
@@
Table XI: Quantitative Evaluation on target probing task. Probes trained for our VisPer-LM perform significantly better as compared to the probes trained on baseline LLaVA-1.5 liu2023improvedllava .